Determine how many bit strings of length 5 can be formed, where the first bit is 0 and three consecutive 0s are not allowed.

1
Expert's answer
2021-12-14T02:54:20-0500

The number of elements of the set of different binary strings of length 5 is "2^5=32."

Determine the number of bit strings of length 5 that contain three consecutive 0s.




"\begin{matrix}
 0 & 0 & 0 & 0 & 0\\
 0 & 0 & 0 & 0 & 1\\
 0 & 0 & 0 & 1 & 0\\
 0 & 0 & 0 & 1 & 1\\
\\
 1 & 0 & 0 & 0 & 0\\
 1 & 0 & 0 & 0 & 1\\
\\
 0 & 1 & 0 & 0 & 0\\
 1 & 1 & 0 & 0 & 0\\
\end{matrix}"

How many bit strings of length 5 can be formed ,where three consecutive 0s are not allowed

32−8=24